OpenStudy (anonymous):

Use 1.7 as the value of Square root of 3. what is the area of equilateral triangle ABC? ___ sq yds

OpenStudy (anonymous):

Area of an equilateral triangle = 1/2 base x heighth. The heighth is measured from the apex (top) of the triangle to the mid-point of the base. Since it's equilateral, all sides are of the same length. So the mid-point of the base is 4yds. Since one side, 1/2 the base and the heighth form a right-triangle, A^2 + B^2 = C^2 (C being the hypotenuse). So C^2 - A^2 = B^2 (B being the heighth) 8^2 -4^2 = B^2.. Take the square root of B^2 and multiply it by half the base (8 yds/2) and that is your answer

OpenStudy (anonymous):